# Cosmological constant as an integration constant

### v3

@@ Abstract
Discussion — In this article, we considered some theories in which the CC arises as an integration constant, including trace-free Einstein gravity, the Codazzi equation, and conformal Killing gravity. We revisited a recent debate [33, 34, 35, 36] concerning the issue of nonuniqueness of solutions to equations of Cotton gravity in symmetry-reduced settiongs. By taking additional derivatives of the Codazzi equations, we find that solutions of the Codazzi equations must satisfy at least one of two conditions: either the solutions coincide with that of the Einstein equations (with arbitrary CC), or that certain contractions of the curvature tensor, the Codazzi tensor, and their derivatives must vanish (in particular, all terms on the right-hand side of Eq. (24) must vanish). The latter can be satisfied in situations of high symmetry, even if the solutions do not satisfy the Einstein equations, but if we regard such non-Einstein solutions as physical, the right-hand side of Eq. (24) must vanish exactly even under perturbations. It is unlikely that such conditions can be satisfied unless the perturbations are highly restricted, and one may in this manner argue that such non-Einstein solutions are unphysical. We find that the Codazzi equations can in principle be closed with the addition of the divergence-free constraint so that such non-Einstein solutions can be excluded at the level of initial data.
@@
We formulated a rather general variational principle that can yield the trace-free Einstein equations, the Codazzi equation, and the equations of conformal Killing gravity. An action for trace-free gravity without a unimodular constraint [38] was recently proposed (another possible approach is described in [23, *Tapia:1998qt, *Tapia:1998as]), but the matter sector was not explicitly incorporated in their treatment; our action (30) provides an alternative that can readily incorporate the matter sector via the action (32). The appearance of Lagrange multipliers and auxiliary fields in the actions corresponds to introducing additional nondynamical degrees of freedom—one might prefer an action that depends exclusively on dynamical degrees of freedom. Moreover, there is some question of whether auxiliary fields in gravity theories are pathological [39, 40, 41]. Regarding the latter, there may be instances in which some of the pathologies (particularly those associated with sharp density gradients) can be mitigated [42, 43, 44], and one can perhaps regard auxiliary fields as effective descriptions for underlying dynamical degrees of freedom (with large masses). Regarding the former concern, it is well-known that one can express the Lagrangian for $f(R)$ theories in the O’Hanlon [45] form $\psi R-V(\psi)$ (see also [46, 47, 48, 49, 50] and references therein), where $\psi$ is an auxiliary field, and one can easily recover the $f(R)$ Lagrangian from the O’Hanlon Lagrangian by algebraic elimination. Of course, a naive algebraic elimination would trivialize the actions considered here, but the point is that auxiliary fields can provide a starting point for constructing an action exclusively from dynamical degrees of freedom. To clarify, we are not suggesting that the variational principle we have proposed is fundamental—indeed, one can construct a variational principle for virtually any set of consistent field equations in this manner. [^4] Presumably one could attempt to use the same strategy to construct a variational principle for inconsistent field equations, but this will result in an action that does not admit extrema. Instead, we argue that it can serve as a first step towards the construction of a more fundamental action for these theories.
@@
Ghosts are a concern; for instance, the Codazzi equation contains up to three derivatives of the metric. One might also recognize that the duplicate matter degrees of freedom $\chi$ in Eq. (32) are necessarily ghosts. At the classical level, there is no issue in the matter sector, as the matter field equations ultimately reduce to those identical to that obtained in the standard variational principle—there is no runaway behavior (see also [51]). In the gravitational sector, one can place restrictions on the initial data such that the system becomes dynamically equivalent to the Einstein equations with a CC. Regarding the Hamiltonian, we note that on shell, both the matter and gravitational actions vanish in our approach, and it is not too difficult to convince oneself that the total matter Hamiltonian must also vanish on shell. Though there is no issue at the classical level, the appearance of ghosts may be of particular concern for the quantization of the theory. Quantization is beyond the scope of the present article (see the discussion in [16] regarding ghosts in higher curvature theories), but we leave the reader with a couple of remarks. First, as argued in the preceding paragraph, the action presented in this article may require further modification at high energies, where the structure of the resulting action may differ significantly. Second, the vanishing of the Hamiltonian indicates that these theories suffer from the problem of time, in which the Hamiltonian operator does not yield the time evolution of the quantum state of the field, a well-known problem in canonical approaches to quantum gravity [52, *Isham:1992ms].
@@
Acknowledgments — We thank Marco de Cesare, Naresh Dadhich, and Victor Tapia for helpful comments and references. PC and JCF acknowledge support from the Taiwan National Science and Technology Council (NSTC) under project numbers 110-2112-M-002-031 and 112-2811-M-002-132, and also from the Leung Center for Cosmology and Particle Astrophysics (LeCosPA), National Taiwan University.

### v4

@@ 5 Discussion
We revisited a recent debate [15; 51; 16; 52] concerning the issue of nonuniqueness of solutions to equations of Cotton gravity in symmetry-reduced settiongs. By taking additional derivatives of the Codazzi equations, we find that solutions of the Codazzi equations must satisfy at least one of two conditions: either the solutions coincide with that of the Einstein equations (with arbitrary CC), or that certain contractions of the curvature tensor, the Codazzi tensor, and their derivatives must vanish (in particular, all terms on the right-hand side of Eq. (24) must vanish). The latter can be satisfied in situations of high symmetry, even if the solutions do not satisfy the Einstein equations, but if we regard such non-Einstein solutions as physical, the right-hand side of Eq. (24) must vanish exactly even under perturbations. It is unlikely that such conditions can be satisfied unless the perturbations are highly restricted, and one may in this manner argue that such non-Einstein solutions are unphysical. We find that the Codazzi equations can in principle be closed with the addition of the divergence-free constraint so that such non-Einstein solutions can be excluded at the level of initial data.
@@
We formulated a rather general variational principle that can yield the trace-free Einstein equations, the Codazzi equation, and the equations of conformal Killing gravity. An action for trace-free gravity without a unimodular constraint [36] was recently proposed (another possible approach is described in [55; 54; 56]), but the matter sector was not explicitly incorporated in their treatment; our action (30) provides an alternative that can readily incorporate the matter sector via the action (32). The appearance of Lagrange multipliers and auxiliary fields in the actions amounts to introducing additional nondynamical degrees of freedom—one might prefer an action that depends exclusively on dynamical degrees of freedom. Moreover, there is some question of whether auxiliary fields in gravity theories are pathological [44; 45; 57]. Regarding the latter, there may be instances in which some of the pathologies (particularly those associated with sharp density gradients) can be mitigated [33; 8; 25], though it is well-known that one can express the Lagrangian for $f(R)$ theories in the O’Hanlon [40] form $\psi R-V(\psi)$ (see also [48; 19; 50; 38; 39] and references therein), where $\psi$ is an auxiliary field, and one can easily recover the $f(R)$ Lagrangian from the O’Hanlon Lagrangian by algebraic elimination. Of course, a naive algebraic elimination would trivialize the actions considered here, but the point is that auxiliary fields can provide a starting point for constructing an action exclusively from dynamical degrees of freedom.
From a more fundamental perspective, one regard auxiliary fields and Lagrange multipliers as vestiges of heavy dynamical degrees of freedom in a more fundamental theory. From this perspective, the actions we propose are not fundamental, but are effective actions obtained after integrating out heavy dynamical degrees of freedom (see, for instance [2]). In this manner, we are in principle trading an finely tuned cancellation for a scenario with additional heavy degrees of freedom—this is not an unreasonable stance, since from quantum gravity considerations, one expects new physics at scales approaching the Planck scale. In this manner, one may view our variational principle as a bottom-up [^5] An interesting question for future investigation is whether such a variational strategy is compatible with the complementary top-down approach of the Swampland program [7; 43]. strategy for constructing theories in which certain parameters (the CC in our case) are entirely insensitive to short-distance physics. While the variational principle we have proposed is not fundamental—indeed, one can construct a variational principle for virtually any [^6] Presumably one could attempt to use the same strategy to construct a variational principle for inconsistent field equations, but this will result in an action that does not admit extrema. set of consistent field equations in this manner—it may provide a useful first step towards the construction of a more fundamental action for these theories.
@@
Ghosts are a concern; for instance, the Codazzi equation contains up to three derivatives of the metric. One might also recognize that the duplicate matter degrees of freedom $\chi$ in Eq. (32) are necessarily ghosts. At the classical level, there is no issue in the matter sector, as the matter field equations ultimately reduce to those identical to that obtained in the standard variational principle, and the ghost fields are constrained to be equal to the original matter fields—there is no runaway behavior (see also [20] for another way to avoid runaway behavior). In the gravitational sector, one can place restrictions on the initial data such that the system becomes dynamically equivalent to the Einstein equations with a CC. Regarding the Hamiltonian, we note that on shell, both the matter and gravitational actions vanish in our approach, and it is not too difficult to convince oneself that the total matter Hamiltonian must also vanish on shell. Though there is no issue at the classical level, the appearance of ghosts may be of particular concern for the quantization of the theory. Quantization is beyond the scope of the present article (see the discussion in [14] regarding ghosts in higher curvature theories), but we offer a couple of remarks. First, as argued in the preceding paragraph, the action presented in this article may require further modification at high energies, where the structure of the resulting action may differ significantly. Second, the vanishing of the Hamiltonian indicates that these theories suffer from the problem of time, in which the Hamiltonian operator does not yield the time evolution of the quantum state of the field, a well-known problem in canonical approaches to quantum gravity [3; 30], though we note that in the framework of effective field theory, general relativity as a perturbative quantum field theory remains predictive at low energies [49]. In any case, the further development of the matter sector may be needed, particularly at the quantum level.
While we have shown that these approaches are in principle viable, the actions we have proposed are not austere in the sense that we have made no attempt to propose a minimal set of fields and symmetries (or other fundamental principles) to constrain the action in a unique manner, as is often done in the construction of field theories. Such constraints would be useful for limiting the allowed modifications to these theories, which in principle form a rather large class (in particular, one can modify the theories we have considered adding to Eqs. (1) and (15) for instance any transverse and trace-free tensor $S_{ab}$ that one can construct from the metric and curvature). Of course, we again emphasize that the actions proposed here are not intended to be fundamental, and should be regarded as effective descriptions (formulated at the classical level).
It may be that the approaches we have considered here ultimately trade austerity and completeness (as we do not expect our theories to be fundamental) for the avoidance of the extreme fine-tuning problem arising from vacuum energy contributions to the CC—some readers may regard this as too high a price to pay. On the other hand, there is some hope that perhaps one can reorganize the action and integrate out the appropriate degrees of freedom to obtain a more natural form for the action in which some principle or symmetry may be more easily identified. A fair question, which we have not addressed, concerns the mechanism that sets the value for the CC; while a fine-tuning in the early universe can be avoided so long as the field equations remain insensitive to the cosmological term, a mechanism for determining the initial data specifying the CC is needed. In short, our work shows that the class of approaches we have considered are viable according to some basic criteria, but at the same time, this approach illustrates some of the challenges and trade offs that one must make in order to avoid these problems. In any case, we hope that our work can provide some guidance for future exploration of scenarios in which the CC arises as an integration constant.
Acknowledgements
@@
We thank Marco de Cesare, Naresh Dadhich, and Victor Tapia for helpful comments and references. PC and JCF acknowledge support from the Taiwan National Science and Technology Council (NSTC) under project numbers 110-2112-M-002-031 and 112-2811-M-002-132, and also from the Leung Center for Cosmology and Particle Astrophysics (LeCosPA), National Taiwan University.